OpenStudy (pottersheep):

Derive y= (3x-9)^9(x+2)^6 I get y1= 162(3x-9)^8(x+2)^5?

OpenStudy (pottersheep):

The answer is 45x(3x-9)^8(x+2)^5...I don't get it!

OpenStudy (anonymous):

you must use the product rule combined with using the chain rule when you find each derivative.

OpenStudy (pottersheep):

Yep that's what I did! Should I show you my steps?

OpenStudy (anonymous):

yes please, let me take a look

OpenStudy (pottersheep):

y = (3x-9)^9(x+2)^6 y1 = 9(3)(3x-8)^8 (6)(1)(x+2)^5 y1 = 162 (3x-9)^8(x+2)^5

OpenStudy (pottersheep):

Thank you

OpenStudy (pottersheep):

Have I gone wrong somewhere?

OpenStudy (anonymous):

yes you forgot to put an addition sign. the product formula is a*d/dx(b) + d/dx(a) * b

OpenStudy (pottersheep):

OHHHHHHH

OpenStudy (anonymous):

when you take the d/dx of the a and the b terms, that's when you execute the sign rule. this is an annoying question lol it will take some time to compute it all

OpenStudy (pottersheep):

Thank you!!!!!

OpenStudy (anonymous):

no problem!